### v2

# A Collectivist, Economic Perspective on AI

Abstract
Information technology is in the midst of a revolution in which omnipresent data collection and machine learning are impacting the human world as never before. The word “intelligence” is being used as a North Star for the development of this technology, with human cognition viewed as a baseline. This view neglects the fact that humans are social animals and that much of our intelligence is social and cultural in origin. Moreover, failing to properly situate aspects of intelligence at the social level contributes to the treatment of the societal consequences of technology as an afterthought. The path forward is not merely more data and compute, and not merely more attention paid to cognitive or symbolic representations, but a thorough blending of economic and social concepts with computational and inferential concepts at the level of algorithm design.
1 Introduction
The current dialogue on artificial intelligence (AI)—in the media, in academia, in industry, in government commissions, and around dinner tables—often seems untethered to reality. The discussion tends to focus on whether one should be on the side of hype or hysteria—AI will either solve humanity’s most pressing problems and usher in a new era of plenty or it will destroy or enslave the human species. Whereas previous eras of rapid technological development were accompanied by such dialogue, the current extreme nature of the hype and hysteria seems unprecedented.
The phrase “AI” arose in the 1950s, and while the phrase was provocative and exciting, the action over the ensuing decades in computer science was elsewhere—in the development of hardware, languages, networks, search engines, human-computer interaction, and eventually data collection and machine learning. The phrase “machine learning” (ML) was coined by an AI researcher (Samuel, 1959), but it was eventually adopted by researchers in many other fields, including operations research, control theory, and statistics, who brought with them a range of experience and applications in engineering and science. Machine learning thus served as an intellectual bridge for a data-intensive era—catalyzing the formation of cross-disciplinary connections and (critically) connecting mathematically inclined researchers from various backgrounds with the computer scientists who were building computing systems and networks of ever-increasing scale.
And then came large-language models (LLMs). The ideas and architectures underlying LLMs were fully in the ML tradition—using gradient-based methods to adjust parameters in large-scale predictive systems—with the key novelty that the data was human language, in truly massive quantities. The output of LLMs was (strikingly) fluent language, giving the appearance of a human-like entity. This triggered the return to prominence of the phrase “AI.”
But whereas an LLM may appear to be a single “entity” that is human-like, it is equally well understood as a “collectivist” artifact. Indeed, in interacting with an LLM, one is interacting implicitly with a vast number of humans who have contributed micro-level data, opinions, linguistic constructions, and creative works to the LLM via the Internet and other media. When these human contributions agree in various ways, the LLM is able to promote that agreement into abstractions that are useful and that strengthen the illusion of personhood. But, while the analogy of an LLM to a person seems irresistible, an analogy of an LLM to a culture is equally valid. Cultures are repositories of narratives, opinions, and abstractions. Cultures have personalities.
Moreover, if we wish to view LLMs as exhibiting human-like intelligence, and in particular as possessing real-world problem-solving skills, then we need to remember that human intelligence is in part social in nature, and that the success of individual efforts is often best measured in the overall social context in which they are embedded (even from the point of view of the individual). Accordingly, let us lift our eyes from the LLM to consider the ecosystems in which LLM algorithms are being embedded. Consider in particular the planetary-scale networks that have emerged in our era to solve problems in domains such as commerce, healthcare, transportation, logistics, education and entertainment. These networks involve vast numbers of heterogeneous participants, some of whom are human and some non-human. The participants are linked by flows of data that increasingly allow them to learn from each other. Learning may involve cooperation, competition, association or collusion. Participants may share some of their resources, including global resources, but they may also want to hold some resources locally. This will occur for many reasons, but most significantly because participants will have a desire to obtain value and competitive advantage from their particular knowledge, data, or creative output.
Thus, overall, as I will argue, an appropriate metaphor for emerging AI systems is closer to that of a market than a search engine, a chatbot, or a personal secretary. (The latter are mere roles in the overall market). This economic perspective allows a balanced consideration of both the producer role and the consumer role of the humans participating in the system. The consideration of the producer role has lagged that of the consumer role, and people are increasingly asking what benefits accrue to them when their creative activity is used in the training of LLMs. Such questions are not new—they also arose in the era of the search engine—but in that case, individuals obtained clear value both as producers and consumers. In their role as producers, they obtained visibility and traffic, and in their roles as consumers they obtained access to information, knowledge, and services. These benefits were part of an implicit social and economic contract in which data on the Internet was treated as free for the taking, and in exchange services were provided freely.
That same contract is being offered (implicitly) by the companies developing LLMs. But the scope of LLMs and generative AI goes far beyond search engines—the goal is no longer to merely provide links that guide users to useful websites, but to aggregate and transform their input data so as to engage in sustained dialog and creative activities. A consequence is that the LLM becomes the endpoint rather than an intermediary, and the benefit of visibility and traffic for producers begins to wither—it is no longer part of the implicit contract.
What will these new markets powered by data and machine learning look like? As in the case of historical markets, bottom-up self-organization will be the dominant paradigm for growth of learning-catalyzed markets. But such growth need not be uncontrolled or outside of our comprehension.
The essential point is that these new markets are arising not because of a deep scientific understanding of the nature of human intelligence, but rather because of the flowering of the concept of an algorithm—one of one of the key achievements of the 20th century. Algorithms, as instantiated in hardware, software and mathematical models, are the way that information technology systems are developed. We can gain leverage over emerging systems by considering more deeply the source of the algorithmic ideas that are driving the technology, asking about the extent to which they connect to societal-level goals, and expanding the space of algorithms if they prove deficient.
In considering the nature and source of algorithmic ideas, it is useful to start with the phrase “computational thinking.” This phrase aims to capture the idea that the algorithmic concepts developed in computer science—such as modularity, abstraction, and scaling—have broad applicability to problem-solving activities throughout science and engineering (Wing, 2006). Indeed, it has been precisely these concepts that have led to the ecosystem that has produced the LLM.
But “computational thinking” arose in the design of systems that had limited, carefully designed interaction with the outside world. The real world—the one that humans operate in—brings two major new sets of issues. First, the real world is characterized by vast complexity and partial observability, such that coping with uncertainty becomes a major issue. Secondly, there is a need to interact in social environments consisting of strategic agents. I believe that these issues are of such vast importance that we cannot simply rely on existing computational design principles to address them. Rather, we need to recognize that there are other ways to conceive of algorithm design. I will propose two other thinking styles that complement computational thinking. I refer to these styles—which are also the fruit of decades of experience—as “inferential thinking” and “economic thinking.” The output of such thinking can be specific algorithms (and analyses of algorithms), although the jargon is often different—in the inferential fields algorithms are often referred to as “procedures” and in the economic fields algorithms are often referred to as “mechanisms.” That these algorithms can be embodied in computational devices is what has given them new power, but focusing solely on “computation” misses the point. It is the thinking behind the algorithms that is important.
Figure 1: Three core thinking styles that have come together in pairwise blends as academic disciplines.
Before turning to concrete instances of design using these thinking styles, let’s return to the issue of coping with uncertainty. Rather than attempting to delimit the notion of uncertainty with a precise definition, let’s instead be informal and consider the broad problem of making decisions when some of the information that would be useful in making the decision is not available. Different fields have contributed different kinds of algorithms to cope with uncertainty. Briefly, the field of statistics has focused on uncertainty arising from sampling, whereby the data available for algorithmic processing is a subset of all of the data that would be useful in principle to solve some problem. The field of economics has also made use of sampling ideas, but a distinctive focus of economics has been a different source of uncertainty—the information asymmetry that arises when an agent interacts with another agent who possesses private knowledge and who may strategically reveal or conceal aspects of that knowledge in order obtain to some desired outcome. Information asymmetry does not tend to go away when the sample size grows. Finally, another source of uncertainty comes from the “when, where, and who” of data collection. When data is collected in the past, or at a distant location, or from a different source than the current source, the uncertainty should grow. The field of computer science has contributed tools to address this problem via the algorithmic concept of provenance, whereby the origin and type of data are tracked systematically and coherently. This gives substance to notions of relevance in decision-making that are often treated informally in other fields.
It is useful to repeat this exercise with other properties that intelligent systems that operate in the real world might be hoped to exhibit, including robustness, coping with bias, understanding causality, and exercising an ability to do experiments. The three thinking styles have contributed diverse and complementary perspectives in each of these cases.
In the next few sections I present vignettes that aim to capture the complementary nature of computational thinking, economic thinking, and inferential thinking. I specifically aim to demonstrate how blends of these thinking styles yield powerful ways to approach system design for emerging data-catalyzed systems. As shown in Figure 1, pairwise blends of these thinking styles have already emerged as academic disciplines. But each of the pairwise blends have made only limited use of the third ingredient, and accordingly they address only part of the problem in systems involving people, machines, and data. What is needed is the tripartite blend.
2 Computation and Inference in Database Design
Let us begin by introducing the perspective of “inferential thinking” in a stylized database problem. [^1] See Figure 4 in Appendix A for a figure that provides a visual accompaniment of the following discussion in the form of a set of flow diagrams. Consider a bank that maintains a database in which the rows correspond to clients and the columns correspond to financial data associated with each client. The bank or others (e.g., auditors) may wish to perform various operations on the data, from simple calculations such as finding the average balance across clients, or more elaborate computations involving spotting unusual transactions. The bank may also wish to provide a privacy guarantee to clients in response to such queries and thus may incorporate randomization. It will also be necessary to track provenance, provide interfaces for clients, and provide long-term storage of data. In short, a great deal of computational thinking will need to go into the deployment of such a database.
What might we mean by “inferential thinking” in database design? On the one hand, the database can implement statistical operations, such as computing a standard deviation or a linear regression. Such computations are in the realm of what a statistician would call “descriptive” but they are not necessarily examples of “inferential thinking.” To clarify the distinction, consider a different database in which the rows are patients in a hospital and the columns are vital signs for the patients, as well as indicators of treatment and responses to treatments. Now, at query time, we might like to ask how likely a particular patient is to respond favorably to a particular treatment. In contrast to the banking example, we are probably not interested in patients who were in the original dataset—we’re interested in “new” patients who come from the same population as the original patients. Indeed, the original patients may be dead and gone, but the data is still valuable. “Inferential thinking” refers to the design and analysis of algorithms that can extract this value. It requires consideration of the underlying population, the set of possible queries, and the design of the sampling operator. It involves methods for checking whether the assumptions made in the design are reasonable post hoc. Although the result of such design methodology is a set of algorithms, the thinking behind the design and analysis goes beyond “computational thinking” in its focus on entities that haven’t been seen before. For such entities the goal is not only to make a prediction, but also to provide a measure of uncertainty regarding that prediction.
More generally, inferential thinking involves characterizations of populations via a generative model, attempting to delineate underlying mechanisms by which data might arise, and choosing a model that fits the data well. Such efforts often fall under the topic of causal inference, where a key concept is the “what if” question—what if the database were different in some way from the data we collected? What if a patient had been given the treatment rather than the control? Can estimates of population-level treatment effects be obtained from the sample data? These issues are subtle; see, e.g., Hernán and Robins (2020).
3 Inference and Incentives
Let us now bring economic thinking into the picture. There are many existing connections of economics to computation; most notably, the field of algorithmic game theory (Nisan et al., 2007). Moreover, within economics proper, the fields of mechanism design (Hurwicz and Reiter, 2006) and information design (Bergemann and Morris, 2019) have been centered around algorithm analysis and design for several decades. My goal is to highlight the opportunities that arise when the kinds of algorithmic ideas that one finds in these fields are juxtaposed with inferential and computational thinking, specifically in the design of large-scale, collectivist machine learning system.
Microeconomics focuses on the choices of strategic agents that are pursuing goals that may be entirely personal. There is also a focus on the overall social welfare that can be achieved when self-interested agents interact. Thus the issues that drive algorithm design in economics include information asymmetries, incentives, social goals, and solutions expressed as equilibria rather than optima.
Let us begin by considering incentives, specifically in the context of inferential problems involving data analysis. Such problems arise often in real-world ML deployments—even if they are not often not treated explicitly. They arise in particular when the suppliers of data are agents who have strategic interests in the outcome of data analysis (Perdomo et al., 2020). This may lead to a misalignment between the goals of the agent and the goals of the data analyst, and lead to competition among multiple agents. In such settings, the system designer will need to consider the design of incentives that will shape behavior; in particular, inducing agents to participate truthfully by sending actual data rather than falsified data or sending data that is selected in some strategic way.
The theory of incentives builds on game theory, which is a mathematical description of strategic behavior (Roger B, 1991). Indeed, the field of mechanism design, which encompasses the study of incentives, can be viewed as the inverse of game theory: Whereas game theory aims to predict the outcome when strategic agents interact—with the outcome expressed as various kinds of equilibria (e.g., Nash equilibria for simultaneous play and Stackelberg equilibria for sequential play)—mechanism design starts with a desired outcome and asks what game would deliver that outcome as an equilibrium.
Sequential play is of particular interest for large-scale collectivist systems given that agents are likely to act asynchronously in such systems. Focusing on just two agents, one agent (referred to as a Leader) plays first, and the other agent (referred to as a Follower) plays next, with the Leader anticipating the Follower’s response. The uncertainty that is present in this situation is one that differs from statistical uncertainty. Known as information asymmetry, it reflects the fact that agents know different things and that there are strategic reasons to withhold one’s knowledge in a transaction. This kind of uncertainty does not go away by mere sampling; rather, it requires the design of an economic mechanism.
One mechanism that is appropriate for sequential play is known as a contract (Laffont and Martimort, 2002). Briefly, the idea is that the Leader does not simply play a single action (e.g., offer a price for some good), but rather presents a menu of options to the Follower, which consists of a set of services and prices. The Follower uses their private knowledge to pick the best option for themselves, and if the Leader has designed the menu well, then many Followers will find an appealing option in the menu. Compared to a mechanism in which a single fixed price is chosen, a contract can deliver higher revenue given that some Followers may opt for a relatively high price in exchange for the services being offered (those Followers with a high “willingness-to-pay”). Moreover, contracts can deliver high social welfare (the aggregrate of the difference between willingness-to-pay and the actual amount paid). Success in achieving such a criterion is expressed in terms of an equilibrium concept; namely, the Stackelberg equilibrium. Algorithm design involves specifying mechanisms (typically decentralized) that yield good Stackelberg equilibria.
Classical contract theory does not incorporate a role for inference from data, but an emerging field of “statistical contract theory” does precisely that (Bates et al., 2024). As an example, let us consider a sequential setting in which the Leader is a buyer who wants to perform hypothesis testing—making “buy” and “no buy” decisions for a sequence of products, where the products are supplied by self-interested suppliers (in the role of Followers). The buyer may be viewed as a marketplace and the testing is needed to determine which products go to market. Suppose that some of the products are of high quality and others are of low quality. The buyer does not know which products are of high quality and may therefore collect a certain amount of data (for example, using a focus group) to make their decision. In doing so, they will be be making statistical errors, given limitations on the amount of data collected (which may be costly). Specifically, there will be false positives and false negatives. We imagine that the buyer’s goal is to minimize some function of these errors. To do so, they have to cope with information asymmetry. The suppliers may know which of their proffered products are of high quality and which are of low quality or may have some rough prior information—in particular they may have invested more effort in making some of the products—but they are not incentivized to reveal this information to the buyer. Indeed, their hope is that some of their low-quality products may end as false positives, which is profit for them.
Statistical contract theory aims to design contracts that incentivize the suppliers to send in products that are more likely to be of high quality, such that the overall mix of products has a controlled statistical error. The contract is a menu of options, where each option involves various costs (e.g., shouldering some of the data-collection burden) and various licensing terms (e.g., responsibilities vis-a-vis customers). Some options will be more or less risky and more or less lucrative. The supplier uses their internal knowledge to make the choice, essentially making a bet on the possible outcomes. The buyer sets up the contract so that it is incentive compatible—in particular, if an item is actually of low quality, the expected total profit for the supplier is nonpositive, so the system cannot be gamed.
Bates et al. (2024) prove that statistical contracts are incentive-compatible in this hypothesis-testing problem if and only the options can be expressed as e-values. An e-value is a function of data that is less than or equal to one in expectation if a null hypothesis is true (Ramdas and Wang, 2025). [^2] E-values can viewed as an alternative to p-values, which are tail probabilities under the null hypothesis rather than expectations. They have a betting interpretation as the multiplicative factor by which wealth increases (in expectation) under the null hypothesis. More generally, when data arrives sequentially in time, the appropriate function of data is the nonnegative supermartingale, which is an e-value at any stopping time and which can be viewed as the accumulation of evidence over time. What the result of Bates et al. (2024) shows is that an important inferential concept (e-values for hypothesis testing) is closely linked to an important economic concept (information asymmetry in contract design).
4 Multi-Way and Multi-Layered Markets, the Internet, and Foundation Models
I now turn to a consideration of examples in which the three thinking styles come together in system design. Rather than attempting to provide formal treatments of such blends, I focus on qualitative examples that have real-world significance.
The Internet is a huge collection of text, images, video, and links and, as such, can be viewed as a source for data analysis. But it is also a place where interactions among humans occur, where creative collective activity such as Wikipedia has taken place, and where markets have arisen. Many of these markets have created real social value, but many are also defective along one or more dimensions—in particular, in their inability to reward creators, to value data as an economic good, to create trust, and to disincentivize socially harmful behavior. Part of the problem is that little thought has been given to mechanism design in building out the Internet. An exception is advertising markets, which have created revenue but which have been a mixed bag with respect to social welfare. In this section, I will discuss other markets that exist on the Internet, or could exist, from the point of view of our three thinking styles, highlighting opportunities for improved social welfare.
4.1 Recommendation systems
Recommendation systems are a classical example of ML systems that are collectivist. In one instantiation of a recommendation system, one considers a graph which links customers on one side and products on the other. Purchases are represented by edges between a customer and the products that they purchase, and graph-theoretic methods exploit similarity patterns in the graph to make targeted recommendations.
Although recommendation systems do bring ML closer to microeconomic considerations, they are limited as microeconomic entities—in particular, no money changes hands. Good recommendations may lead customers to make purchases, but conceptually this is just a way to make an existing market for physical goods more efficient. There is no strong need for consideration of incentives.
Let us consider a market that is currently in the midst of technology-driven change—that of recorded music. In the days of yore recorded music was a physical good but it has become a virtual good. For virtual goods, the lack of economic mechanism design in recommendation systems is problematic, leading to an impoverished reward system for creators, who wield little market power. Let us consider an alternative. In Figure 2, I depict the design of a three-way market for recorded music. [^3] This market is the architecture underlying the company United Masters, where the author is a board member and consulted on the market design. At one vertex are musicians, who supply songs, and at a second vertex are listeners. The musicians and listeners are linked by a classical ML-based recommendation system. Critically, there is a third vertex, which are brands. Brands often make use of music in products and outreach and they need well-chosen music which fits their image and connects well with the demographic they cater to. Additional recommendation systems, also powered by trained ML models, provide these connections. Moreover, critically, the overall design incorporates incentives. When a brand needs a song, they are supplied with a song from a particular artist (using an ML model), and the artist is paid in that moment. Audience reaction is measured. Other brands can see that reaction, and if it happens to be aligned with a demographic they’re also interested in, then they are incentized to reach out to the artist and partner with them. [^4] It is worth noting that United Masters has signed over 1.5 million musicians to date, and their music is used by brands such as the NBA, Bose, and State Farm. This is an instance of a collectivist AI system that has created jobs.
Figure 2: A depiction of a three-way market for recorded music which links musicians, listeners and brands.
Note the difference with the classic online business model for recorded music, where musicians upload their music to the cloud and it is streamed to listeners for free. Money is made by the platform, via subscriptions or by advertising, but there is no direct connection between producer and consumer, and there is a weak incentive for the platform to send money back to the musicians. [^5] Indeed, there’s a strong incentive for the platform to use generative AI tools to replace the musicians!
Different ways of conceiving of market dynamics for a learning-based online service can have rather different outcomes in terms of social welfare.
4.2 Data markets
Let us now consider a different problem domain that features a three-component market, in this case taking the form of a layered structure (Fallah et al., 2024). In Figure 3, a user is shown interacting with a platform, which provides a service and receives a payment in return (concretely, let us suppose that the platforms provide access to credit for a fee). We imagine that the platform can learn from the data that it obtains from the user and thereby improve the service.
Thus far we have a market where data plays an informational role, but data is not a transacted good. Both the user and the platform are incentivized to engage in this market, depending on the details of the service and the fee. But in many such situations, the platform does not make enough revenue from the fees it collects to realize a profit (Cross, 2023). Thus, it turns to a set of third-party data buyers. These buyers wish to acquire data for their own purposes (such as carrying out market research). The platform acts as a supplier, and the data becomes a transacted good. It will be priced according to its value to the data buyer and other factors. Again, the incentives are aligned for the platform and the buyer.
But now consider the overall system and ask whether the user is incentivized to participate in this layered market. A new issue has arisen—the user has lost control over their privacy in this market. Whereas before the platform could be held to a contract so that the user has a guarantee that their private data is being used in a limited way—and they are receiving a service and therefore are presumably willing to incur some privacy loss—now the user is told that third-party data buyers are acquiring his or her data and all bets are off. No new service accrues to the user in exchange. The privacy loss can be unbounded and substantial.
Thus, the user is likely to walk away, and the design needs to be elaborated for the market to function. Let us imagine, for example, that platforms decide to provide a formal guarantee of privacy when sending data to third-party data buyers. Concretely, this would mean that noise is added to the data at a level that is contractually specified (and can be audited). Although the noise level could be subject to government regulation, let’s instead leave the choice in the hands of the platforms. A platform is incentivized to provide a nontrivial level of noise, because users will shop among platforms to find one that provides a desirable level of privacy in conjunction with an effective service. Moreover, such a platform, by accruing more users, will collect more data and can make further improvements to its service. On the other hand, data buyers are averse to noise and will presumably pay less for data from the platforms that provide a stronger privacy guarantee. Thus, there is a conflict for the platforms. The way to understand how the conflict will play out is to model the overall system as a game (it is a generalized Stackelberg game) and find its equilibria. This requires specifying utility functions for the various players.
Figure 3: A three-layer market in which a platform provides services, and also sells data to third-party buyers.
In this scenario, both the platforms and the data buyers will likely be ML systems, learning from the data that they receive. Moreover, data is an endogeneous part of an overall system in which learning, data, and human preferences all interact. Understanding how the overall system will behave—and in particular whether it will work at all—requires a collectivist perspective that combines ML with economics.
4.3 Foundation models, bias, and local knowledge
Foundation models are large-scale ML systems that aim to make high-quality predictions in domains of major scientific or societal interest. For example, LLMs are foundation models for natural language and AlphaFold is a foundation model for protein structure (et al., 2019).
The phrase “high-quality predictions” warrants some discussion. Let us consider first scientific domains, where there is in principle a ground truth to compare to. For example, AlphaFold exhibits high overall accuracy when compared to protein structures that have been determined in the lab. This does not mean, however, that it is uniformly accurate. Rather, it is accurate on training data and test data that have been obtained from past scientific investigation. Unfortunately, scientists are often interested in phenomena that are on the edge of knowledge, such that there may be little past data to support accurate prediction. Indeed, Angelopoulos et al. (2023) showed that AlphaFold can given highly biased confidence intervals (intervals that are overly narrow and do not cover the ground truth) for certain queries involving proteins that exhibit quantum fluctuations (where there are not many ground-truth measurements). Angelopoulos et al. (2023) also demonstrated that such biased confidence intervals arise in a wide range of other scientific domains.
This problem can be addressed via a technique developed by Angelopoulos et al. (2023) that is known as prediction-powered inference. This is an inferential algorithm in which assessments of uncertainty obtained from global foundation models are adjusted based on local ground-truth measurements that are possessed by a local agent. Such local knowledge may involve measurements that were not available in the training of a foundation model or may simply reflect a particular (desirable) bias of the local agent. PPI-corrected confidence intervals provably cover ground-truth estimands, under standard statistical assumptions.
More generally, in domains involving strategic interactions between humans and non-humans, an agent that queries another agent (which may be a foundation model) will generally need to be concerned about bias. Moreover, the strategic nature of the interaction means that the bias may have been created willfully, to align with the goals of the other agent. In this context, PPI can be viewed as something more than a debiasing technique. If the agent providing the data or the foundation model is aware that the receiving agent will be making use of local ground-truth data, then that agent will be disincentivized from providing data that is significantly biased. The agent will also be incentivized to expand the scope of their data or their model to meet the receiving agent’s needs (if they want to continue to interact with that agent).
Still more generally, ambiguity in what constitutes the “correct response” will arise from the fact that data and knowledge are often local, contextual, and fleeting. Solving one’s local problem will often involve a blend of information from outside sources with information that is only available locally in space and in time.
5 Discussion
The study of multiple agents in computer science is by no means new, and researchers in fields such as multi-agent ML, human-computer interaction, and algorithmic game theory will recognize that it is their work that is being promoted here, as are the perspectives of the social sciences broadly speaking. Moreover, there are antecedents of our arguments in the study of “collective intelligence” (Tumer and Wolpert, 2004; Parkes, 2004; Malone and Bernstein, 2015). That literature has two main branches, one focused on qualitative, experimental work with groups, and the other focused on artificial collectives with designed agent utilities. My focus is different. I want to uncover algorithmic design principles for emerging real-world ML-based systems in which many of the participants are human and many are non-human. The goals and utilities of the humans are to be understood and respected, not designed. The principles should be simultaneously computational, economic, and inferential.
Although ML-powered markets can be expected to differ in important ways from classical markets, it’s worthwhile to recall some of the appealing features of classical markets. First, they provide a form of uncertainty reduction—buyers can count on products to be available and plan accordingly. Second, they cope with heterogeneity. Third, they create new roles on an as-needed basis. It is early days for ML-powered markets, but it is not hard to imagine many new roles that can be expected to arise, akin to those that have arisen in previous eras of technology but based around data and learning—auditors, brokers, aggregators, sellers, buyers, artists, forecasters, insurers, and explorers. These roles will give rise to personalized services, economies of scale, and appropriate touch points at which regulatory control can be exerted to mitigate those biases that are proscribed by legal or ethical considerations.
Design principles based on our tripartite blend make it easier to discuss issues of interest to both individuals and collectives, such as privacy, fairness, ownership, alignment, reputation, and transparency. The blend allows these issues to be treated as tradeoffs rather than being reduced to black-and-white distinctions. For example, differential privacy is a computational method that adds noise to data to guarantee privacy according to a particular definition (Dwork and Roth, 2014). But this is only part of the story. An individual’s decision to ask for a certain level of differentiable privacy in an interaction will trade off with the other costs and benefits that the individual expects from the interaction. Moreover, inferential issues arise whenever noise is being added to data, and there are quantitative tradeoffs that can be formulated that link inferential accuracy and privacy (Duchi et al., 2014). Making full use of the tripartite blend will generally involve solutions that take the form of tradeoffs.
Finally, for AI to grow into a mature engineering discipline that delivers systems that yield value to humans in real-world settings, it will need far more than blends of existing algorithms (and far more than just more data and more compute). In this regard, it’s useful to learn from the history of the fields of chemical engineering and electrical engineering, which brought the complex phenomena of chemical reactions and electromagnetism under control. This was achieved by developing modular, transparent design concepts that were appropriate for the phenomena. The modularity allowed large systems to be designed piecemeal, allowed system failures to be diagnosed and repaired, and allowed multiple stakeholders to participate in the evolution and regulation of systems. We are far from such design concepts in the current stage of development of AI. Moreover, chemical engineering and electrical engineering had at their foundations Schrödinger’s equation and Maxwell’s equations, solid foundations that could guide the development of simplifying modular approximations in the face of exceedingly complex phenomena.
For AI, we certainly have exceedingly complex phenomena—cognitive, social, commercial, and scientific—but we do not have the equivalent of Maxwell’s equations as a guide. We are winging it. Going forward, we therefore need the very best of our overarching, hard-won general scientific and humanistic principles—including rationality, experimentation, dialog, openness, cooperation, skepticism, empathy, and humility—as daily companions on the journey ahead.
Appendix A Inference and Privacy in Databases
We discussed three database design scenarios in Section 2. Figure 4 provides a visual depiction of these scenarios.
In Figure 4(a), we depict the standard privacy-preservation problem, in which the problem is to develop a (randomized) algorithm $Q$ that provides privacy guarantees while ensuring that the privatized response $\tilde{y}$ to a query $x$ is not too far from the “true response” $y$. The latter refers to the response that would be obtained without privacy constraints.
Note that the diagram is not meant to suggest that the randomized algorithm $Q$, which concretely could be a differentially private operator, needs to involve transforming the entire database. Rather, using standard ideas in the differential privacy and online learning literatures (cf. Dwork and Roth, 2014), we envision $Q$ as being called on an as-needed, possibly query-dependent basis.
Figure 4(b) presents an inferential perspective, in which the database is assumed to arise from an underlying population, under a (randomized) sampling operator $S$. The sampling operator may be given or designed (or it may be some blend). The goal is to ensure that $y$ is not too far from the “true response” $y^{*}$. Here the latter is an unobservable population quantity. Statistical theory is needed to make this kind of guarantee.
Figure 4: Database design problems that highlight the complementary nature of computational thinking and inferential thinking.
Of course, in real-world problems there is often a need to combine these two thinking styles. A more general setup is one in which the goal is predictive inference for a new patient, but where we also want to protect the privacy of the individuals in the original dataset, perhaps because their offspring will have an interest in ensuring their own privacy. Figure 4(c) presents the privacy-preserving-inference problem, where we wish to make predictions for new individuals who are sampled from the underlying population while providing a privacy guarantee for the individuals who were in the original database. Here, we want $\tilde{y}$ to be close to $y^{*}$ (see, e.g., Blum et al., 2008; Duchi et al., 2014).
Privacy appears in several of our vignettes, including the database example and the three-layer data market. This is not surprising, given its prominence in public discussion of information technology, including in legal circles, and given that differential privacy and other computational procedures have provided ways to quantify certain aspects of privacy. As we’ve noted, this opens up a discussion of tradeoffs involving economic value and inferential value.
It’s also worth pointing to another computational framework for treating issues surrounding privacy that is known as contextual integrity (Nissenbaum, 2004). Contextual integrity aims to provide formal reasoning strategies for evaluating data flows in social contexts. This is a different direction than the quantitative and algorithmic computation-economics-inference focus of the current paper, but the goals of connecting social desiderata to technological design are shared, as is the emphasis on context. It is of significant interest to engage in a deeper comparison of the current perspective with that of contextual integrity, not only for privacy but for related economic and personal desiderata in AI systems.
Finally, there is much to discuss regarding the interactions of privacy with incentives and trust. This is a domain in which hardware and cryptography can provide some assistance—if an individual is assured that his or her data is being processed using some form of trusted execution environment and/or secure multi-party communication algorithm (Liao et al., 2025), then that individual can be viewed as incentivized to tolerate additional privacy risk.
Appendix B The Missing Middle Kingdom in AI Education
What are the implications of the tripartite blend for undergraduate and graduate education? A naive response would be that students should simply take a collection of classes in computer science, economics, and statistics. But this response is lacking in several ways. First, these three disciplines have not cornered the market on “computational thinking,” “economic thinking” and “inferential thinking.” Important instances of these thinking styles can be found throughout academia. Second, classes in computer science, economics, and statistics often aim to solve problems from previous eras, and the tripartite blend aims at emerging problems. Third, by simply asking students to take more classes we are putting the onus on the students to put together the concepts they are learning in the context of real-world problems. Educators should be shouldering more of the burden of innovative thinking in an era of major change.
It is noteworthy that the phrase “computational thinking” arose in part to support an expanded presence of computational ideas in academic curricula (Wing, 2006). In 2015 my colleagues and I at UC Berkeley built on that perspective to design a new course, Data 8, that aimed to introduce college freshmen to a blending of “computational thinking” and “inferential thinking.” To give a concrete example of what this blend can mean at the freshman level, consider a problem in which the goal is decide whether a treatment is effective based on a comparison of two groups of outcomes. Students were presented with two columns of numbers, labeled “treatment” and “control,” and asked whether it seemed plausible these columns were significantly different (i.e., there was an effect of the treatment), or not. To aid this assessment, students first learned how to make histograms (learning just enough Python programming to carry out this task). Comparing the histograms visually gave a sense of possible differences. To take a step further and address an inferential question, we introduced the permutation test. This test involves some “what-if” thinking, where one considers the possibility that the columns of numbers come from the same source. In that world, one can lump the two columns together, permute the resulting single column, split it into two new columns, and compute whatever comparative statistic between the columns that one likes. (We used the variation distance, which is a simple comparison between histograms). Doing this repeatedly yields a sampling distribution, which can be plotted as a histogram. Placing the statistic of the original two columns on this histogram, one can assess whether this statistic seems likely given the hypothetical “what-if” thinking that the simulation process embodies. Finding that statistic out in the tail of the histogram suggests that that hypothetical assumption may be false.
Clearly there is a lot of inferential reasoning going on in this example, and it can be blended with computational concepts in ways that go beyond just plotting histograms. For example, how does one permute a column of numbers in Python? There is a naive $O(n^{2})$ algorithm to do it, but there is also a more sophisticated $O(n)$ algorithm. Aha! Then, critically, with this blend in hand, we returned to the students and asked them to find their own problems, and accompanying datasets, where “treatment” and “control” were whatever real-world distinction they want to investigate. (E.g., does water quality vary between more wealthy parts of a city and well wealthy parts? or Was there more deforestation in the Amazon this past year as compared to five years ago?). Such real-world investigation was exciting for a diverse set of students, in ways that the problems studied in classical computer science curricula are sometimes not (e.g., game-playing programs). Data 8 became the fastest-growing class in the history of UC Berkeley, and is currently taken by over 1500 students each semester.
Returning to the original design process for Data 8, it is natural to consider adding “economic thinking” to the mix. As I have emphasized in this article, in many problems data arises out of interactions among humans, and the algorithms that analyze such data should take such interactions into account. One can well imagine that basic ideas such as matching markets, contract design, and auctions could be integrated into the algorithms-for-real-world-problem-solving perspective of Data 8.
Such material would retain the core focus on algorithms, but it would also open up coverage of social issues (such as privacy, fairness, ownership, alignment, reputation, and transparency). It would also demonstrate that the world of algorithm design—the key leverage point that we have on the evolution of modern technology—is worth understanding by a broad swath of students, not just computer engineers. Understanding the role of algorithms that address economic and inferential issues can help engender a dialog that is more multi-directional than simply teaching computational concepts to constituencies outside of computer science.
A related virtue of the tripartite blend is that it is inclusive of the wide range of disciplines that have already developed connections to computer science, economics, and statistics. Indeed, from these three starting points, one arrives quickly at disciplines such as cognitive science, social psychology, law, sociology, public policy, mathematics, physics, biology, and the humanities. These are all disciplines that can provide additional perspectives that are essential in shaping human-centric technology. To take but one example where these perspectives link to our discussion, consider behavioral economics, which links cognitive and social psychology with economics, aiming to overcome a perceived overreliance on unrealistic assumptions of rationality in classical economics. Here too, however, existing concepts need to be revisited in the data-centric era. Recalling that humans provide the data on which AI artifacts such as LLMs are based, we see that behavioral data is in some sense already baked into learning-based systems. Blending the massive-but-messy information in such data with the more traditional, experiment-driven insights coming from behavioral economics is an important agenda item for future research.
It is often suggested that there should be stronger links between computer science and the humanities. But in practice this often boils down to having computer science students take an ethics class or another elective, and to having humanities students acquire programming skills. The gap between computer science and the humanities is large, and it is difficult to design direct bridges. This is due in part to the absence of connections that flow through the social sciences in attempts to design such bridges. Indeed, some social sciences have natural connections to computer science (e.g., economics), others have natural connections to the humanities (e.g., history), and in general the focus on humans and groups of humans provides essential perspectives. Our tripartite blend can be viewed as providing a new kind of engineering core that aims to catalyze existing connections and spawn others. It provides a middle kingdom between engineering and the humanities that can help refresh academic debates and help design forward-looking, inclusive curricula.
Appendix C LLMs, Uncertainty, and Collectives
I have argued that the ability to cope with uncertainty—at both the individual level and the collective level—is a hallmark of intelligence. Recall that we are considering a broad, informal notion of uncertainty that involves sampling, information asymmetry, provenance, and other ways in which information that is relevant to a decision may be biased, partially available, or stale.
Current-generation LLMs make little use of any explicit notion of uncertainty. They can give the impression that they are quantifying uncertainty and reasoning inductively, but this is in part an illusion, arising from the fact that the humans who produced the data on which the learning algorithms are based employed terminology and arguments associated with reasoning under uncertainty. It is far from clear that LLMs have extracted useful general principles for managing uncertainty from such data, particularly given that individual humans do not necessarily follow normative rules (see, e.g., Gilovich et al., 2002; Griffiths, 2024). Moreover, some of the known problems with LLMs, such as systematic overconfidence (Sun et al., 2025), are not exhibited by individual humans to the degree that they are exhibited by LLMs. It remains an important item for the research agenda to develop a deeper understanding of the cognitive science of uncertainty and the appropriate decision-theoretic underpinnings for LLMs.
But, as I have argued, a cognitivist perspective is not sufficient for understanding and coping with real-world uncertainty. Moreover, I have suggested that LLMs can be usefully viewed as a collectivist entity rather than a singular intelligence. In this section I provide further discussion of these two issues and their interrelationships.
A first point is that collectives play an essential role in changing the world in ways that shape and mitigate uncertainty. An individual human foraging for fruit may or may not succeed in finding fruit on any given day, but once a collective creates a market for produce, the uncertainty regarding finding fruit drops significantly. A human can then depend on fruit being available and build on that certainty, perhaps opening a pastry shop.
Collectives also help to define which uncertainties an individual should focus on and how uncertainties impact the choice of actions. Consider an example from behavioral psychology—a two-choice maze experiment in which a mouse learns that there is more food on the left branch than on the right branch in a ratio of 2/1. The mouse now needs to decide which branch to take on the next trial. A mouse that is decision-theoretically optimal would visit the left branch with probability one, maximizing its chances of finding food. Real mice (and humans in analogous situations) often do something different, known as “probability matching”—they visit the left branch twice as often as the right branch.
Why should that be? Consider a setting in which there are many mice. Here, if each mouse goes left with probability one, then the food on the right is an unexploited resource. A probability-matching algorithm rectifies this—each mouse independently chooses to go left versus right with a ratio of 2/1 and the result is high overall social welfare. An appealing explanation for its emergence is that it is an equilibrium that is found via evolutionary forces (Brennan and Lo, 2011).
In general, collectives are the appropriate level of analysis for many problems involving scarcity and uncertainty. The solutions to such problems may be baked in by evolution, but more often than not they require communication and interaction among agents. This leads us to microeconomic concepts but also requires us to augment those concepts with a fuller toolbox for the management of uncertainty (for example, the distribution-free methods such as conformal inference that have been developed in recent years (Angelopoulos and Bates, 2023)). Thus we should envisage new kinds of markets in which data flows and distributed statistical analysis play key roles in shaping agent behavior. Machine learning concepts will be crucial given the kind of dynamic, highly distributed systems that are our likely future, in which data will be increasingly partial, noisy, biased, local, contextual, strategic, and generally incoherent. From this perspective, intelligence is not just about knowing things, and giving answers that are “correct,” but knowing how to act when one’s knowledge is partial, and knowing how to interact with other individuals whose knowledge is partial.
In summary, the two themes of “social” and “uncertainty” are closely related. [^6] It’s worth noting that this is a point that has been emphasized in the field of finance (Merton, 1995). On the one hand, social environments create various kinds of uncertainty, including information asymmetry (some actors are better informed than others, in ways that depend on the problem at hand) and strategic obfuscation (am I being told the truth and how would I know?). On the other hand, social environments permit cooperation and the sharing of information, mitigating uncertainty for everyone, and thereby improving decision-making and enabling long-term planning. Real-world intelligence is as much a social, communications, economic, and cultural concept as a cognitive concept.